A ball thrown up from a rooftop of height 54m lands on the ground in 5s. What is its speed 13m below?


1
Expert's answer
2021-02-11T08:53:18-0500

Let's first find the initial speed of the ball from the kinematic equation:


"y=y_0+v_0t-\dfrac{1}{2}gt^2,""0=54+5v_0-122.5,""v_0=13.7\ \dfrac{m}{s}."

Let's find the time that the ball takes to reach the height 13 meters below rooftop:


"41=54+13.7t-4.9t^2,""4.9t^2-13.7t-13=0."

This quadratic equation has two roots: "t_1=3.54\ s" and "t_2=-0.75\ s". Since time can't be negative, the correct answer is "t=3.54\ s".

Finally, we can find the speed of the ball 13 meters below rooftop:


"v=v_0-gt,""v=13.7\ \dfrac{m}{s}-9.8\ \dfrac{m}{s^2}\cdot3.54\ s=-21\ \dfrac{m}{s}."

The sign minus means that the speed of the ball directed downward.